Elaborated derivation of de equation


         y=Asin(wt-θ)


where w=omega=2π/T


       θ=phase difference between the particles

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear venu,

The distance of the particle from mean position in a particular direction at anyinstant of time is known as instantaneous displacement.It is given by

Y = A sin (ωt + φ)If the particle starts from Mean position, φ = 0 theni) Y = 0, at Mean positionii) Y = A at extreme positioniii) Y =2A at θ = ωt = 30° = π/6 or after a time interval t = T/12 sec.iv) Y =2A at θ = ωt = 45° = π/4 or t = T/8 sec.v) Y =23 A at θ = ωt = 60° =π/3 or t =T/6 sec.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi